bank provides two different kind of bank account: current accounts and savings accounts. Every bank customer has one or both of these. 90% of bank customers have current accounts and 60% of bank customers have savings accounts. If a customer is chosen uniformly at random from the set of all bank customers, calculate the following probabilities


 

a) The probability they have a current account or a savings account.

 

b) The probability they have a current account and a savings account.

 

c) The probability they have a current account but not a savings account.

 

The probability they do not have a savings account, given that they have a current account


1
Expert's answer
2021-10-26T10:30:04-0400

Solution:

Denote the persons having Savings account by S

and persons having Current account by C.

"P(S)=60\%=0.6
\\P(C)=90\%=0.9"

(a):

Every bank customer has one or both of these. It means

The probability they have a current account or a savings account"=P(S\cup C)=100\%=1"

(b):

b) The probability they have a current account and a savings account"=P(S\cap C)"

"=P(S)+P(C)-P(S\cup C)
\\=0.6+0.9-1
\\=0.5
\\=50\%"

(c): The probability they have a current account but not a savings account"=P(C \cap S')"

"=P(C)-P(C \cap S)
\\=0.9-0.5
\\=0.4
\\=40\%"

(d):

The probability they do not have a savings account, given that they have a current account"=P(S'|C)
\\=\dfrac{P(S'\cap C)}{P(C)}
\\=\dfrac{0.4}{0.9}
\\=\dfrac49"